OpenStudy (anonymous):

1. Simplify: (3 + 5) × 10 - 10

OpenStudy (anonymous):

18 43 0 70

OpenStudy (anonymous):

70

OpenStudy (anonymous):

3+5 = 8 8*10 = 80 80-10= 70 (:

OpenStudy (anonymous):

2. Simplify. 3 × 22 + 8 ÷ 2 - (5.6 + 3.7)

OpenStudy (anonymous):

60.7

OpenStudy (anonymous):

Simplify: 50 - 3 × (6 - 4)3

OpenStudy (anonymous):

32 68 26 74

OpenStudy (anonymous):

is that suppose to be an exponent at the end ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

26

OpenStudy (anonymous):

how many questions do you have lol?

OpenStudy (anonymous):

5. Simplify: 19 - (2 + 8) - 20 ÷ 5

OpenStudy (anonymous):

10 but i will anwser the rest

OpenStudy (anonymous):

this one is 5

OpenStudy (anonymous):

if you need help on any other ones just let me know (:

OpenStudy (anonymous):

kay`